OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2?

OpenStudy (anonymous):

2x + y = 4 2x - y = 4 x + 2y = 4 x - 2y = 4

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

use the point slope form of a line \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

y+3= 1/2(x+2)??

OpenStudy (anonymous):

@freckles